# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

### v1

@@ 15 Runtime and Computational Complexity
17 Sampling approach from network features
Our approach revolves around the strategic sampling of edges, emphasizing their importance in the overall network structure. We introduce innovative method for this purpose, such as one leveraging edge betweenness centrality (EBC). In these proposed methods, our approach to edge sampling is guided by the assessed significance of edges with respect to the topological structure of the graph. This methodology extends random sampling techniques commonly employed in contemporary graph embedding, which typically rely on information derived solely from sparse graph connectivity, encompassing nodes within a 1-hop distance on the graph. To elaborate on our sampling approach, we extend the definition of the network $G=(V,\mathbf{W})$ to a triplet $G=(V,\mathbf{W},\gamma)$ such that $\gamma:V\rightarrow(0,1]$ is some probability distribution over $V$ (or $\mathbf{W}$) providing the significance of each node/edge. We choose to estimate $\gamma$ directly from the graph network using kernel density estimation (KDE).
The kernel density estimator of the distribution ${\gamma}({{V})}$, is given by
$\hat{\gamma}(x)=\frac{1}{Nh}\sum_{i\in V}{\frac{\mathbf{K}(\bar{v}(i)-x)}{h}}$ (29)
Where $\bar{v}$ accumulate statistics measuring the extent of diffusion spread using SGW among nodes, $\mathbf{K}$ is symmetric kernel function (chosen to be a smooth Gaussian symmetric function), $h>0$ is the smoothing bandwidth that controls the amount of smoothing.
In a similar way, one can also construct a probability measure $\gamma:V\times V\rightarrow(0,1]$ describing the significant of an edge. In this case, we choose to calculate the distribution from the edge betweenness centrality measurements (EBC), $\left\{w_{i,j}^{BE}|i,j\in V,i\sim j\right\}$ (see Eq. 31) where we use kernel density estimation (KDE) to estimate the density distribution. Let $\gamma:V\times V\rightarrow\mathbb{R}^{+}$ denote the probability distribution over edges in the graph $G=(V,\mathbf{W})$, using the distribution of EBC, (EBC is defined in Eq. 30). We denote the estimated distribution calculated directly from the EBC measurements as $\gamma({\mathbf{W^{BE}})}$. Formally, the kernel density estimator of distribution of EBC, ${\gamma}({\mathbf{W^{BE}})}$ is given by
$\hat{\gamma}(x)=\frac{1}{N(e)h}\sum_{i\sim j}{\frac{\mathbf{K}(w_{i,j}^{BE}-x)}{h}}$ (30)
Where $N(e)$ is a fixed number of positive edges to be sampled.
The advantage of using KDE is that it is a non-parametric density estimator, which does not require assumption that the underlying density function is from a parametric family. Note that KDE works by estimating the density at each point (node or edge) as a weighted sum of the densities of neighboring points, with the weights given by a kernel function $\mathbf{K}$. In our experiments, we used the Gaussian kernel for simplicity, but one can use other kernel choices.
Figure 4: Sampling from Betweenness Centrality distribution
17.1 Sample from Network Features using Edge Betweenness Centrality
Motivation: We employ Edge Betweenness Centrality (EBC) based sampling to create a strategy that samples edges from the EBC distribution. This enables us to optimize embeddings by predominantly using edges with low EBC, which are concentrated in dense clusters. In contrast, high-EBC edges typically occur in transitions between clusters and form a sparser set that is less frequently sampled. During optimization with SGD and negative sampling, these high-EBC edges play a crucial role in revealing the global structure of the network and connections between different clusters.
Consider the illustration in Fig. 4, where edges marked with black arrows correspond to edges with high edge betweenness centrality values, indicating their importance in connecting different dense clusters. Sampling edges with low EBC assists in identifying dominant local structures, while infrequently sampling high betweenness centrality edges is critical for understanding large clusters and the network’s global structure.
  Input: $G=(V,\mathbf{W})$, kernel density function $\mathbf{K}$, bandwidth parameter $h$, $N_{e}$ for the number of edge samples.
  Step 1: Construct the centrality graph $G=(V,\mathbf{W}^{BE})$ using Eq. (31).
  Step 2: Utilize $\mathbf{K}$, $\bar{v}$, and $\mathbf{W}^{BE}$ in Kernel Density Estimation (Eq. 30) to estimate $\hat{\gamma}$.
1  Step 3: Initialize a list of positive edges $C_{e}=\left\{\right\}$. Set $i=1$.
  Step 4: While $i\leq N_{e}$ do:
  Step 5: Sample an edge $w_{ij}$ using $\hat{\gamma}$ , $w_{ij}\sim\hat{\gamma}(\mathbf{W}^{BE})$.
  Step 6: Add $w_{ij}$ to $C_{e}$, and and set $i=i+1$.
  Step 7: Repeat steps 4-6. STOP when $i=N_{e}$
  Output: Set of edges $C_{e}$ for positive samples in Algorithm 1.
Algorithm 2 Edge Betweenness Centrality Sampling
The Edge Betweenness Centrality:
Edge Betweenness Centrality (EBC) quantifies edge importance based on the fraction of shortest paths that pass through it. For a graph with vertices and edge weights, the EBC of an edge $(i,j)$ is given by:
$w_{i,j}^{BE}=\sum_{k\neq t}\frac{\sigma_{kt}(w_{ij})}{\sigma_{kt}}$ (31)
Here, $\sigma_{kt}$ is the number of shortest paths from node $k$ to node $t$, and $\sigma_{kt}(w_{ij})$ is the number of those paths that include the edge $w_{ij}$. The centrality graph $G=(V,\mathbf{W}^{BE})$ is constructed using these EBC values, modifying the edges of the original graph $G$.
Sampling Strategy of EBC using KDE estimator:
Let $\gamma:V\times V\rightarrow\mathbb{R}^{+}$ denote the probability distribution over edges in the graph $G=(V,\mathbf{W})$, using the distribution of EBC. We denote the estimated distribution calculated directly from the EBC measurements as $\gamma({\mathbf{W^{BE}})}$.
Given the set of edge betweenness centrality $\left\{w_{i,j}^{BE}|i,j\in V,i\sim j\right\}$ as an input to the Kernel Density Estimation (KDE) estimator, the density function is estimated automatically. With a fixed number of positive edges $N(e)$ to be sampled, we sample $N(e)$ edges from the estimated distribution. Although we used the Gaussian kernel in our experiments, other kernel choices are viable. In EBC based-sampling, we sample from the distribution of edge betweenness centrality measurements. This approach prioritizes the sampling of edges based on their importance in the graph structure. By treating edges as bottlenecks, we sample high betweenness centrality edges less frequently, thereby revealing clusters. Edge betweenness centrality serves as a measure of an edge’s bottleneck potential, guiding our SGD sampling and prioritizing edges based on their relevance and significance in capturing underlying relationships and connectivity patterns in the graph. Combined with the proposed encoding and initial embedding of multi-scale graph structure, the approach provides a comprehensive representation of the data.

### v4

@@ 16 Runtime and Computational Complexity
Remark 3: Optimal Choice of coefficients $a_{k}$ for Minimizing $\Lambda_{\psi}$ for optimal choice of $\Lambda_{\psi}$ in the multi-scale setting.
To find the coefficients $a_{t_{j},k}$ that minimize $\Lambda_{\psi}$ where:
$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}$ (20)
subject to the constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$, we optimize the coefficients $a_{t_{j},k}$, which control how much of the signal at scale $t_{j}$ and power $k$ is captured by the SGW operator. We use the Lagrangian $\mathcal{F}$, incorporating the energy constraint into the optimization problem and Define:
$\mathcal{F}(a_{t_{1},0},\dots,a_{t_{J},K},\mu)=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}+\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}-1\right)$ (21)
[…]
Taking partial derivatives, solving for $a_{t_{j},k}$, and differentiating $\mathcal{L}$ with respect to each coefficient $a_{t_{j},k}$:
$\frac{\partial\mathcal{F}}{\partial a_{t_{j},k}}=\frac{-a_{t_{j},k}/\Lambda^{2k}}{\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}}+2\mu a_{t_{j},k}=0.$ (22)
rearranging the equation above we obtain:
$\frac{a_{t_{j},k}}{\Lambda^{2k}}=2\mu a_{t_{j},k}\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$
assuming $a_{t_{j},k}\neq 0$, we can divide both sides of the equation by $a_{t_{j},k}$, yielding:
$\frac{1}{\Lambda^{2k}}=2\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$ (23)
thus the sum $\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}$ is proportional to the inverse of $\Lambda^{2k}$, implying that the optimal wavelet coefficients should decay proportionally to the inverse powers of the Laplacian eigenvalues $a_{t_{j},k}\propto\frac{1}{\Lambda^{k}}$.
Next, applying the energy constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$ and substituting $a_{t_{j},k}=\frac{1}{\Lambda^{k}}$, we obtain:
$\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}=1$ (24)
Thus, the normalized optimal coefficients $a_{t_{j},k}$ are:
$a_{t_{j},k}=\frac{\frac{1}{\Lambda^{k}}}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}}}.$ (25)
Note that the wavelet energy is distributed across all scales and powers of the Laplacian, with a decay that prioritizes low-frequency components (small $k$).
Using the form of the coefficients $a_{t_{j},k}$ above, the expression for $\Lambda_{\psi}$ is:
$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{4k}}}}$ (26)
which is the minimized value of $\Lambda_{\psi}$, showing that the optimal choice of coefficients reduces the overall constant, improving the localization and efficiency of the SGW operator. Thus, the optimal choice of wavelet coefficients $a_{t_{j},k}$ that minimizes $\Lambda_{\psi}$ in the multi-scale SGW operator is proportional to $\frac{1}{\Lambda^{k}}$, prioritizing low-frequency components.